OpenStudy (anonymous):

Can someone give me the answer please. The table below shows two equations: Equation 1 |4x − 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution, and equation 2 has solutions x = −4, 1. The solutions to equation 1 are x = 3, −1.5, and equation 2 has no solution. The solutions to equation 1 are x = 3, −1.5, and equation 2 has solutions x = −4, 1.

OpenStudy (phi):

For the 2nd equation, add -8 to both sides |2x + 3| + 8-8 = 3-8 and simplify. 8-8 is 0 and 3-8 is -5 |2x + 3| = -5 Now, what do you know about | | ? Do you know that it makes whatever is "inside" positive? In other words, the left side can *never be negative* and that means no matter what "x" value you try, you will never get the left side to equal -5